OpenStudy (anonymous):

Someone please help , how do you solve this ? The function f(x) = 1.50x + 12.50 relates how much Kelsey pays for digital service, f(x), to the number of movies, x, she streams in a month. What is the value and meaning of f(18)? $27.00; this is the amount Kelsey pays for 18 months of service $27.00; this is the amount Kelsey pays for 18 movies in 1 month $39.50; this is the amount Kelsey pays for 18 months of service $39.50; this is the amount Kelsey pays for 18 movies in 1 month

OpenStudy (welshfella):

the x intercept is the value of x when y = 0 and y intecept is value of y when x = 0 so to find the x intercept solve for x :- 3x - 2(0) = 18 3x = 18 x = ?

OpenStudy (anonymous):

Hey sorry that was the wrong question because it dosnt make sense it's ... The function f(x) = 1.50x + 12.50 relates how much Kelsey pays for digital service, f(x), to the number of movies, x, she streams in a month. What is the value and meaning of f(18)?

TheSmartOne (thesmartone):

That's that same question though...?

OpenStudy (anonymous):

Oh gosh nvm , and x=15

TheSmartOne (thesmartone):

wrong. You are multiplying the 3 to x. So you have to divide to get rid of it: \(\sf\Large \frac{3x}{3}=\frac{18}{3}\)

OpenStudy (anonymous):

x=6 then ?

TheSmartOne (thesmartone):

correct

OpenStudy (anonymous):

Okay but now what ? What do i do with the 6?

TheSmartOne (thesmartone):

I don't know why @welshfella had you calculate the x-intercept... It is of no importance in the question :/

TheSmartOne (thesmartone):

f(x) = 1.50x + 12.50 They want to know what f(18) = ? so plug in x=18 and tell me what you get :)

OpenStudy (anonymous):

39.5

TheSmartOne (thesmartone):

correct.

TheSmartOne (thesmartone):

so what 2 answer choices can we eliminate?

OpenStudy (anonymous):

A & b

TheSmartOne (thesmartone):

good job and I think welshfella 's comment was for your previous question which you editted :p

OpenStudy (anonymous):

Yes it was haha

TheSmartOne (thesmartone):

read this line: The function f(x) = 1.50x + 12.50 relates how much Kelsey pays for digital service, f(x), to the number of movies, x, she streams in a month. what do you think would be the answer between C and D?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

correct, great job :)